On January 1, 2008, Peace Corporation bought machinery under a contract that required down-payment of P50,000, plus 24 monthly payments of P25,000 each, for a total cash payments of P650,000. The cash equivalent price of the machinery was P550,000. The machinery has an estimated useful life of 10 years. Peace uses the straight-line method of depreciation. How much should Peace report in its 20078 income statement as depreciation for the machinery?
1
Expert's answer
2021-10-29T06:10:02-0400

Answer

The service life is ten years and the fair market value of $110,000 will be applied in calculating the depreciation. The total cash payments of $130,000 since it includes interest as well as the principal. The fair market value of $110,000 is used as the cost price of the machinery. The formula to calculate the depreciation will be as follows:

Annual Depreciation expense = "\frac{Cost\; of\; the\; asset - Salvage\; value}{Useful\; life}"


"\frac{110,000 - 5,000}{10} = \frac{105,000}{10} = 10,500"


Annual depreciation expense = $10,500